OpenStudy (anonymous):

Determine whether each equation has the given ordered pair as a solution. y = 5x + 8; (0, 8)

jimthompson5910 (jim_thompson5910):

Replace x with 0. Replace y with 8

jimthompson5910 (jim_thompson5910):

Then simplify as much as possible

OpenStudy (anonymous):

my problem is i dont know how yet

jimthompson5910 (jim_thompson5910):

y = 5x + 8 8 = 5(0) + 8 8 = 0 + 8 8 = 8 So (0,8) is a solution

OpenStudy (anonymous):

lol thanks you make it sound so easy

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

i wish i could figure it all out lol

jimthompson5910 (jim_thompson5910):

you'll get it with enough practice

jimthompson5910 (jim_thompson5910):

just keep at it

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

sure thing